OpenStudy (anonymous):

Compute the derivative of g(t) if g(t)=sqrt(t^3+4t+1)

OpenStudy (anonymous):

\[\frac{d}{dx}\sqrt{f(x)}=\frac{f'(x)}{2\sqrt{f(x)}}\] by the all mighty chain rule

OpenStudy (anonymous):

\[g'(t)=\frac{ 1 }{ 2(t^3+4t+1)^1/2 (3t^2+4) }\]